OpenStudy (anonymous):

The points P( 3,5) and Q (7, -1) are endpoints of a diameter of a circle. What is the lenght of a radius of the circle?

OpenStudy (anonymous):

do you know how to find the distance between 2 points?

OpenStudy (callisto):

First, find the length of diameter of the circle using distance formula. \[distance = \sqrt{(x_2 - x_1)^2 + (y_2-y_1)^2}\]Put \((x_1 , y_1) = (3,5) \) and \((x_2 , y_2) = (7,-1) \) into the formula. \[diameter = \sqrt{(7 - 3)^2 + ((-1)-5)^2}= \sqrt{4^2 + (-6)^2}= ...\] Radius = diameter /2